OpenStudy (anonymous):

which expression is equivalent to -(5x+4)? A) -9x B) 9x C) 5x-4 D) -5x-4

OpenStudy (steve816):

Distribute!

OpenStudy (steve816):

Think of the - as just -1.

OpenStudy (zzr0ck3r):

\(-(5x+4)=-1(5x+4)\) and \(a(b+c)=ab+ac\)

OpenStudy (steve816):

Do you think you can distribute -1(5x+4) ?

OpenStudy (anonymous):

D?

OpenStudy (steve816):

Yup! You're correct :)

OpenStudy (anonymous):

It was actually C; but thanks for the help anyway. c: @steve816